OpenStudy (anonymous):

Solve for x: 3x^2=450

OpenStudy (anonymous):

first divide both sides by 3, what do you get?

OpenStudy (anonymous):

x^2=150

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now take the square root of both sides, what do you get?

OpenStudy (saifoo.khan):

first divide 450 by 3 x^2 = 450/3 x^2 = 150 Now take sqrt on both sides x = root150 x = 5root6

OpenStudy (anonymous):

Thank you!

OpenStudy (saifoo.khan):

Uw,